# Paterson compactifications, inverse limits and shadowing for Deaconu–Renault systems

### 07 Apr 2026 v1

@@ 2 Metric on infinite product spaces
Observe that $Q$ is surjective. We equip $W_{0}$ with the quotient topology induced from $A$ via the map $Q$, that is, a set $U\subseteq W_{0}$ is open if and only if $Q^{-1}(U)\subseteq A$ is open. Since $Q$ is continuous and $A$ is compact, it follows that $W_{0}$ is compact. Define the length map $\ell:W_{0}\to\mathbb{N}\cup\{\infty\}$ by $\ell(x)=n$ if $x\in X^{n}$ for some $n\geq 0$ and $\ell(x)=\infty$ if $x\in X^{\mathbb{N}}$. Note that $\vec{0}$ is the unique element of $W_{0}$ with length 0.
@@
which is open in $A$ since each $B_{i}$ is open in $X_{\infty}$ and $X_{\infty}\setminus K$ is open in $X_{\infty}$ (because $K$ is compact $X$).
@@
which is open in $A$ as well. Therefore, both $Z[B_{1}\times\cdots\times B_{k},K]$ and $C[K]$ are open in $W_{0}$.
@@
meaning the sequence (in $X$) consisting of the $i$-th coordinates of all those $x^{n}$ whose length is at least $i$.
@@ 3 The inverse limit and its shifts
    If $x\in\widetilde{X}$ and $\ell(x)\geq 1$, then its tail $x_{2}x_{3}\cdots\in\widetilde{X}$. More precisely, if $x=x_{1}x_{2}\ldots\in D_{\infty}$ then $x_{2}x_{3}\cdots\in D_{\infty}$. If $x=x_{1}x_{2}\ldots\in D_{\mathrm{lim}}$ has finite length at least $2$, then $x_{2}x_{3}\cdots\in D_{\mathrm{lim}}$, and if $x$ has length $1$ then $\vec{0}\in D_{\mathrm{lim}}$.
@@
The choice $\operatorname{Dom}(\sigma)=\{x\in\widetilde{X}:\ell(x)\geq 1\}$ is important because it preserves the point $\vec{0}$ in the image, which will be needed later. However, $\sigma$ is not the appropriate inverse of $\alpha_{f}$ at words of length $1$. For this reason we introduce the truncated shift $\widehat{\sigma}=\sigma|_{\{x\in\widetilde{X}:\ell(x)\geq 2\}},$ and it is $\widehat{\sigma}$, rather than $\sigma$, that is inverse to $\alpha_{f}$.
@@
    Moreover, $\vec{0}\in E$ iff $\widetilde{X}$ contains a word of length $1$.
@@ 5 Shadowing via inverse limits for ultrametric Deaconu–Renault systems
Claim. $(y_{i})_{i\geq 0}$ is a $\mathcal{V}_{l}$–pseudo-orbit for $\alpha_{f}$, i.e. $d(\alpha_{f}(y_{i}),y_{i+1})<2^{-l}$ for all $i$.
@@ 5.1 Examples
In this section we present examples of spaces and Deaconu–Renault systems satisfying the hypothesis of Section 2, Section 5, and of Theorem 5.5.
@@
We conclude the paper with two examples of Deaconu–Renault systems satisfying the hypotheses of Theorem 5.5.
@@
The one-point compactification of $D=X\setminus\{p\}$ is canonically homeomorphic to $X$ by adding back the point $p$. Hence we identify $D^{\infty}:=X$, $d_{\infty}:=d,$ so $d_{\infty}$ is a compatible ultrametric on $D^{\infty}$ and its restriction to $D$ is the metric on $D$. Moreover, $f(D)=X=D^{\infty}$ (indeed, for any $y\in X$ we have $0y\in W_{0}\subseteq D$ and $f(0y)=y$), so $D\subseteq f(D)\subseteq D^{\infty}$ as required in Lemma 5.4(i).
@@
Let $(E,\alpha_{f})$ be the Deaconu–Renault system associated to $(X,f)$, as in Section 3. We claim that $\vec{0}\in E$, so that Theorem 5.5 applies in this example. Let $\widetilde{X}$ be as in Section 3. By Proposition 3.11, it suffices to show that $\widetilde{X}$ contains an element of length $1$.
@@
so $\tau$ is finite on $D$ and $D$ is open in $A$ (hence locally compact) but $D\neq X$.
@@
and on each $W_{k}$ we have $\tau\equiv k+1$, so $f|_{W_{k}}=\sigma^{k+1}|_{W_{k}}\colon W_{k}\to A$ is a homeomorphism onto $A$.
@@
Since $D=A\setminus\{0\,1^{\infty}\}$, its one-point compactification $D^{\infty}$ is naturally identified with $A$ (by adding back the missing point $0\,1^{\infty}$). Under this identification we take $D^{\infty}:=A$, $d_{\infty}:=d|_{A},$ which is a compatible ultrametric on $D^{\infty}$; its restriction to $D$ is the metric on $D$ used below. Moreover, $f(D)=A=D^{\infty}$, so $D\subseteq f(D)\subseteq D^{\infty}$ as required in Lemma 5.4(i).
@@
As in the previous example, we show that Theorem 5.5 applies in this case as well. Let $(E,\alpha_{f})$ be the Deaconu–Renault system associated to $(X,f)$, as in Section 3. By Proposition 3.11, it suffices to show that $\widetilde{X}$ contains an element of length $1$.

### 13 Apr 2026 v2

@@ 2 Metric on infinite product spaces
Observe that $Q$ is surjective. We equip $W_{0}$ with the quotient topology induced from AA via the map $Q$, that is, a set $U\subseteq W_{0}$ is open if and only if $Q^{-1}(U)\subseteq A$ is open. Since $Q$ is continuous and AA is compact, it follows that $W_{0}$ is compact. Define the length map $\ell:W_{0}\to\mathbb{N}\cup\{\infty\}$ by $\ell(x)=n$ if $x\in X^{n}$ for some $n\geq 0$ and $\ell(x)=\infty$ if $x\in X^{\mathbb{N}}$. Note that $\vec{0}$ is the unique element of $W_{0}$ with length 0.
@@
which is open in AA since each $B_{i}$ is open in $X_{\infty}$ and $X_{\infty}\setminus K$ is open in $X_{\infty}$ (because $K$ is compact $X$).
@@
which is open in AA as well. Therefore, both $Z[B_{1}\times\cdots\times B_{k},K]$ and $C[K]$ are open in $W_{0}$.
@@
meaning the sequence (in $X$) consisting of the $i$-th coordinates of all those $x^{n}$ whose length is at least ii.
@@ 3 The inverse limit and its shifts
    If $x\in\widetilde{X}$ and $\ell(x)\geq 1$, then its tail $x_{2}x_{3}\cdots\in\widetilde{X}$. More precisely, if $x=x_{1}x_{2}\ldots\in D_{\infty}$ then $x_{2}x_{3}\cdots\in D_{\infty}$. If $x=x_{1}x_{2}\ldots\in D_{\mathrm{lim}}$ has finite length at least 22, then $x_{2}x_{3}\cdots\in D_{\mathrm{lim}}$, and if $x$ has length 11 then $\vec{0}\in D_{\mathrm{lim}}$.
@@
The choice $\operatorname{Dom}(\sigma)=\{x\in\widetilde{X}:\ell(x)\geq 1\}$ is important because it preserves the point $\vec{0}$ in the image, which will be needed later. However, $\sigma$ is not the appropriate inverse of $\alpha_{f}$ at words of length 11. For this reason we introduce the truncated shift $\widehat{\sigma}=\sigma|_{\{x\in\widetilde{X}:\ell(x)\geq 2\}},$ and it is $\widehat{\sigma}$, rather than $\sigma$, that is inverse to $\alpha_{f}$.
@@
    Moreover, $\vec{0}\in E$ iff $\widetilde{X}$ contains a word of length 11.
@@ 5 Shadowing via inverse limits for ultrametric Deaconu–Renault systems
Claim. $(y_{i})_{i\geq 0}$ is a $\mathcal{V}_{l}$–pseudo-orbit for $\alpha_{f}$, i.e. $d(\alpha_{f}(y_{i}),y_{i+1})<2^{-l}$ for all ii.
@@ 5.1 Examples
In this section we present examples of spaces and Deaconu–Renault systems satisfying the hypothesis of Section 2, Section 5, and of Theorem $5.$
@@
We conclude the paper with two examples of Deaconu–Renault systems satisfying the hypotheses of Theorem $5.$
@@
The one-point compactification of $D=X\setminus\{p\}$ is canonically homeomorphic to $X$ by adding back the point pp. Hence we identify $D^{\infty}:=X$, $d_{\infty}:=d,$ so $d_{\infty}$ is a compatible ultrametric on $D^{\infty}$ and its restriction to $D$ is the metric on $D$. Moreover, $f(D)=X=D^{\infty}$ (indeed, for any $y\in X$ we have $0y\in W_{0}\subseteq D$ and $f(0y)=y$), so $D\subseteq f(D)\subseteq D^{\infty}$ as required in Lemma 5.4(i).
@@
Let $(E,\alpha_{f})$ be the Deaconu–Renault system associated to $(X,f)$, as in Section 3. We claim that $\vec{0}\in E$, so that Theorem 5.5 applies in this example. Let $\widetilde{X}$ be as in Section 3. By Proposition 3.11, it suffices to show that $\widetilde{X}$ contains an element of length 11.
@@
so $\tau$ is finite on $D$ and $D$ is open in AA (hence locally compact) but $D\neq X$.
@@
and on each $W_{k}$ we have $\tau\equiv k+1$, so $f|_{W_{k}}=\sigma^{k+1}|_{W_{k}}\colon W_{k}\to A$ is a homeomorphism onto AA.
@@
Since $D=A\setminus\{0\,1^{\infty}\}$, its one-point compactification $D^{\infty}$ is naturally identified with AA (by adding back the missing point $0\,1^{\infty}$). Under this identification we take $D^{\infty}:=A$, $d_{\infty}:=d|_{A},$ which is a compatible ultrametric on $D^{\infty}$; its restriction to $D$ is the metric on $D$ used below. Moreover, $f(D)=A=D^{\infty}$, so $D\subseteq f(D)\subseteq D^{\infty}$ as required in Lemma 5.4(i).
@@
As in the previous example, we show that Theorem 5.5 applies in this case as well. Let $(E,\alpha_{f})$ be the Deaconu–Renault system associated to $(X,f)$, as in Section 3. By Proposition 3.11, it suffices to show that $\widetilde{X}$ contains an element of length 11.